OpenStudy (anonymous):

multiply and simplify expression (10-z)z

OpenStudy (anonymous):

10z-z^2

OpenStudy (skullpatrol):

Do you know the distributive property?

OpenStudy (skullpatrol):

Distributive Property of Multiplication with Respect to Subtraction For all real numbers a, b, and c: $$a(b-c)=ab-ac$$and $$(b-c)a=ba-ca$$

OpenStudy (anonymous):

Yes, I completely forgot and just needed a refresher on this, but thank you

OpenStudy (skullpatrol):

np :)